# Range Membership Inference Attacks

### v1

@@ Appendix E Extra results
Table 3: True Positive Rate under different attacks on different datasets at small false positive rates of $1\%$ and $0.1\%$. MIAs cannot be conducted on incomplete data, so we fill the missing columns with the modes and run the attack on them. Standard deviation over random sampling iterations are reported, except for CelebA, where we use all available candidates.
@@
Table 4: Improvement in AUCs after switching from MIA to RaMIA across multiplue iterations of random sampling. We do not random sample but use all available attack images in CelebA.
Purchase-100 CIFAR-10 CelebA AG News
$\Delta\text{AUC}$ $2.62\pm 0.04$ $4.12\pm 0.06$ $5.4$ $1.20\pm 0.2$
Figure 6: RaMIA gets better when the range size increases on CelebA, even if we assume the attacker cannot sample any training image.
@@
Figure 7: Correlation between the percentile of RaMIA and MIA scores of members among non-members. The larger the percentile is, the more non-members the member dominates, and the more likely for it to be classified as IN. The Pearson correlation coefficients are provided for each plot. It shows that the vulnerability to MIA is correlated with that to RaMIA.

### v2

@@ 5.5 RaMIAs quantify privacy risks more comprehensively than MIAs
Table 4: True Positive Rate under different attacks on different datasets at small false positive rates of $1\%$ and $0.1\%$. MIAs cannot be conducted on incomplete data, so we fill the missing columns with the modes and run the attack on them. Standard deviation over random sampling iterations is reported, except for CelebA, where we use all available candidates. The TPR and FPR are calculated based on the range membership information, as described in Sec 5.4. For a fair comparison, we should compare RaMIA and MIA based on the same membership testing backbone, e.g. MIA with RMIA versus RaMIA with RMIA.
@@
5.6 Factors affecting RaMIA performance
Training data density in the range
(b) RaMIA performs better on CIFAR-10 when the range size increases. The range size here is equal to the number of distinct transformations applied to images.
@@ 5.6 Factors affecting RaMIA performance
Figure 5: Correlation between the percentile of RaMIA and MIA scores of members among non-members. The larger the percentile is, the more non-members the member dominates, and the more likely for it to be classified as IN. The Pearson correlation coefficients are provided for each plot. It shows that the vulnerability to MIA is positively correlated with that to RaMIA.
Susceptibility to MIAs and RaMIAs is correlated
Ranges containing training points that are susceptible to MIAs are also more susceptible to RaMIAs. Researchers have previously discovered that machine learning models memorize duplicate data more [26, 6]. In our CelebA dataset, each celebrity has a different number of photos in the training set, which can be thought of as each identity having different levels of duplication in the training set. Similar to the insights from MIAs, we also observe that identities that have more training images, i.e. higher duplication rate, are more susceptible to RaMIA. Figure 6 shows the relationship between the percentile of each range’s RaMIA score within non-members’ RaMIA scores and the duplication rate. Generally speaking, identities that have more training photos are more prone to RaMIAs. Similarly, correlation can be observed on the other three datasets in our experiments, where the training points’ RaMIA score percentiles among non-members are positively correlated with their MIA score percentiles 5.
@@
5.7 Mismatched training and attack data hurts attack performance
Difference to existing augmentation-based MIAs
Existing attacks [5, 48] also use augmented queries in the attack, but with a different rationale and assumption of the attacker’s knowledge. Since they adopt the existing privacy notion based on point queries, only (augmented) images seen by the model in the training stage are considered as members. Hence, the attacker needs to know the exact train-time augmentations and augment images accordingly to not violate the privacy notion. In RaMIA, the set of augmentations is given by the challenger (Def 2), which can contain augmentations not used in training, but considered as privacy leaking. Using the aggregation method in [5] will hurt the attack performance if non-training augmentations are used. However, RaMIA is designed to be robust in this scenario (Fig 3(c)) due to its trimming process.
5.8 RaMIA on redacted data
Many large language models (LLMs) are trained with sensitive textual data. Some of the data with sensitive information redacted might be publicly available. Similar to our experiment with data with missing values, we can apply RaMIA to redacted data to identify which of them are used to train a target LLM. Accurately identifying the redacted sentences paves the way for reconstructing them as a follow-up attack. Figure 7 shows the results. In this experiment, we use spaCy [18] to mask peoples’ names to simulate the masking of personally identifiable information (PII). We then generate 10 possible sentences for each masked sentence using BERT and conduct RaMIA. The MIA performance is the average attack performance over all 10 possible sentences. The performance gap is smaller compared to that in Figure 3(d). The reason might be that BERT fails to produce diverse PII completions, making all candidate sentences similar to each other, and reducing the power of RaMIA.
Figure 7: RaMIA on a subset of AG News where names are redacted.
6 Conclusion
In this paper, we argue that membership inference attacks are only useful as a privacy audit tool when querying exact copies of training and test data. Moving the query to similar points causes a drastic decrease in performance, rendering MIAs less useful. We conclude MIAs fail to comprehensively capture the notion of privacy, and thus propose a new class of inference attack, RaMIA, that extends the notion of MIAs and covers the failure cases of MIAs by checking if a given range contains a training point. We introduce RaMIA as an attack framework that can be implemented with any existing MIA algorithm. We show that it can provide better privacy auditing with very few samples taken randomly. We hope our work can make more privacy researchers and practitioners aware of the shortcomings of MIAs, and shift their attention to RaMIAs. As it is the first paper that brings up this new framework, there is room for improvement in specific attack algorithms. For example, a better sampling process will surely increase the gap between RaMIA and MIA. Nevertheless, we have shown our framework is sensible and useful. In future work, we hope to design more powerful RaMIA strategies that are robust to the change of membership signals and datasets, especially on LLMs where we believe our privacy notion is extremely relevant.
Appendix A Attack algorithms
LOSS
LOSS [46] computes loss values as a proxy of membership score on given points: $\text{MIA}(x;\theta)=\mathit{l}(x;\theta)$. To compute the likelihood, an easy way is to take the exponential of the negative of the loss $\text{I\kern-1.49994ptP}=\exp^{-\mathit{l}}$.
RMIA
RMIA [48] computes membership score by applying chain rule: $\text{I\kern-1.49994ptP}{(\theta|x)}=\frac{\text{I\kern-1.04996ptP}{(x|\theta)}\text{I\kern-1.04996ptP}{(\theta)}}{\text{I\kern-1.04996ptP}{(x)}}$. The score is then compared with all available population data points to obtain the percentage of population points being dominated by the given point: $\text{I\kern-1.49994ptP}_{z\in Z}{[\frac{\text{I\kern-1.04996ptP}{(\theta|x)}}{\text{I\kern-1.04996ptP}{(\theta|z)}}\geq\gamma]}$, where the term $\text{I\kern-1.49994ptP}{(\theta)}$ will cancel out with each other. The normalizing constant $\text{I\kern-1.49994ptP}{P(x)}$ is computed with reference models: $\text{I\kern-1.49994ptP}{(x)}=0.5\mathbb{E}_{\theta_{\text{IN}}}\text{I\kern-1.49994ptP}{(x|\theta_{\text{IN}})}+0.5\mathbb{E}_{\theta_{\text{OUT}}}\text{I\kern-1.49994ptP}{(x|\theta_{\text{OUT}})}$. In its offline version, the in models are unavailable. In this case, the former probabilities are approximated by the latter term $\text{I\kern-1.49994ptP}_{\text{IN}}=a\text{I\kern-1.49994ptP}_{\text{OUT}}+(1-a)$. The hyperparameter $\alpha$ is chosen based on the reference models. Specifically, one reference model is chosen as the temporary target model, and the rest are used to attack it. The value of $\alpha$ is chosen to be the best-performing value under this setting, obtained via a simple sweeping. In our experiment, we use the offline attack only. The $\alpha$ values for Purchase-100 and CIFAR-10 are taken from [48]. For CelebA, we set it to be 0.33. For AG News, we set it to be 1.0.